### v1

# Global existence and boundedness in an attraction-repulsion chemotaxis system with nonlocal logistic source and sublinear productions

Abstract
This paper deals with the following attraction-repulsion chemotaxis system with nonlocal logistic source and sublinear productions
$\left\{\begin{array}[]{rrll}&&{u}_{t}=d_{1}\Delta{u}-\chi\nabla\cdot({u}^{k}{\nabla{v}})+\xi\nabla\cdot({u}^{k}{\nabla{w}})+\displaystyle{\mu{u}^{m}\left(1-\int_{\Omega}{u}(x,t)\mathrm{d}x\right)},&x\in\Omega,\,t>0,\\ &&{v}_{t}=d_{2}\Delta{v}-\alpha{v}+f(u),&x\in\Omega,\,t>0,\\ &&{w}_{t}=d_{3}\Delta{w}-\beta{w}+f(u),&x\in\Omega,\,t>0,\\ &&\frac{\partial{u}}{\partial\nu}=\frac{\partial{v}}{\partial\nu}=\frac{\partial{w}}{\partial\nu}=0,&x\in\partial\Omega,\,t>0,\\ &&{u}(x,0)={u}_{0},\quad{v}(x,0)={v}_{0},\quad{w}(x,0)={w}_{0},&x\in\Omega,\end{array}\right.$
in an open, bounded domain $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ with smooth boundary $\partial\Omega$. Assume the parameters $d_{1}$, $d_{2}$, $d_{3}$, $\chi$, $\xi$, $\alpha$, $\beta$ and $\mu$ are positive constants, initial data $({u}_{0},{v}_{0},{w}_{0})$ are nonnegative and the function $f(u)\leq Ku^{l}\in C^{1}([0,\infty))$ for some $K,l>0$. Under appropriate conditions on the parameter $k$, $l$ and $m$ we show that the above problem admits a unique globally bounded classical solution. Further, to illustrate the analytical results, some of numerical examples are provided.
keywords:
Attraction-repulsion model, Chemotaxis system, Classical solution, Global existence, Non-local source, Numerical simulation
1 Introduction and motivation
Many researchers are currently involved in the analysis of nonlinear partial differential equations considering the nonlocal source. In recent years, works on dynamical systems with nonlocal source are gaining much attention in mathematical biology [2, 3]. In any biological phenomenon, in general, the cells interact not only through attraction, there exist cases of repulsion between them as well. This necessitates the study of attraction-repulsion mechanism in the context of chemotaxis. Microglia are neuronal support cells in the central nervous system and they interact with secretory signals through chemotaxis to secrete attractive and repulsive chemicals. In the case of Alzheimer’s disease (a type of neurodegenerative disease), the motion of microglias is affected by such chemoattractants and chemorepellents produced by them, which stimulate its aggregation in the central nervous system. This serves as a good motivation to study chemotaxis system involving one species with two chemicals. The process was modelled using the attraction-repulsion Keller-Segel system as in [33] and mathematical studies of similar models have been widely explored thereafter.
This study inherits the following initial-boundary value problem
$\left\{\begin{array}[]{rrll}&&{u}_{t}=d_{1}\Delta{u}-\chi\nabla\cdot({u}^{k}{\nabla{v}})+\xi\nabla\cdot({u}^{k}{\nabla{w}})+\displaystyle{\mu{u}^{m}\left(1-\int_{\Omega}{u}(x,t)\mathrm{d}x\right)},&x\in\Omega,\,t>0,\\ &&{v}_{t}=d_{2}\Delta{v}-\alpha{v}+f(u),&x\in\Omega,\,t>0,\\ &&{w}_{t}=d_{3}\Delta{w}-\beta{w}+f(u),&x\in\Omega,\,t>0,\\ &&\frac{\partial{u}}{\partial\nu}=\frac{\partial{v}}{\partial\nu}=\frac{\partial{w}}{\partial\nu}=0,&x\in\partial\Omega,\,t>0,\\ &&{u}(x,0)={u}_{0},\quad{v}(x,0)={v}_{0},\quad{w}(x,0)={w}_{0},&x\in\Omega,\end{array}\right.$ (1.6)
in an open, bounded domain $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ with smooth boundary $\partial\Omega$ and here $\nu$ denotes the outward unit normal on $\partial\Omega$. We assume the parameters $d_{1}$, $d_{2}$, $d_{3}$, $\chi$, $\xi$, $\alpha$, $\beta$ and $\mu$ are positive constants and the initial data $({u}_{0},{v}_{0},{w}_{0})\in{{C^{0}}(\overline{\Omega})}\times{{W^{1,q}}(\overline{\Omega})}\times{{W^{1,q}}(\overline{\Omega})}$ are nonnegative. The unknown functions ${u}$, ${v}$ and ${w}$ represent the cell density, the concentration of attractive and repulsive chemical substances, respectively. The system (1.6) describes the interaction of one cell with two different chemicals secreted by the cells. The attraction effect is given by the term $-\chi\nabla\cdot(u^{k}\nabla v)$, where $\chi$ is the attraction coefficient. On the other hand, the term $\xi\nabla\cdot(u^{k}\nabla w)$ represent the repulsion effect, where $\xi$ is the repulsion coefficient. The proliferation rate of cells are assumed to have a logistic growth in the form of nonlocal source and is given by $\mu u^{m}(1-\int_{\Omega}u)$, where $\mu$ is the growth rate constant. Here, the decay rates of the chemicals are given by $\alpha v$ and $\beta w$ respectively. Here the growth rate of the cells is named by the function $f(u)$. Further, $d_{1}$, $d_{2}$ and $d_{3}$ are positive diffusion coefficients of cell density and chemical signals respectively.
In addition, we also assume that the initial data ${u}_{0}$, ${v}_{0}$ and ${w}_{0}$ satisfy
$\left\{\begin{array}[]{rrll}&&{u}_{0}\in{\mathcal{C}^{0}(\overline{\Omega})},\quad\mbox{with}\quad{u}_{0}\geq 0\quad\mbox{in}\,\,\Omega,\\ &&{v}_{0},{w}_{0}\in{\mathcal{W}^{1,q}(\overline{\Omega})},\quad\mbox{for some}\,\,{q}>n,\quad\mbox{with}\quad{v}_{0},{w}_{0}\geq 0\quad\mbox{in}\,\,\Omega.\end{array}\right.$ (1.9)
Moreover, the function $f$ is defined as follows
$f(s)\in C^{1}([0,\infty))\quad\mbox{and}\quad 0\leq f(s)\leq Ks^{l}\quad\mbox{for some}\quad K,l>0\quad\mbox{and all}\quad s\geq 0.$ (1.10)
The system (1.6) is the generalization of the mimimal chemotaxis one species and one stimuli system by Keller-Segel in 1970.
$\left\{\begin{array}[]{rrll}{u}_{t}&=&\Delta{u}-\nabla\cdot({u}\nabla{v}),\\ {v}_{t}&=&\Delta{v}-{v}+{u},\end{array}\right.$ (1.13)
where ${u}$ and ${v}$ denote the cell density and the concentration of chemical signals, respectively. Analytic and numerical simulations of various Keller-Segel models and Keller-Segel models with logistic source have been studied in the past few decades. For the recent evolution in this field one can refer the articles by Lankeit and Winkler [24], Bellomo et al. [1] , Horstmann [17] and the references therein.
To the better understanding of the system (1.6), Luca [33] propsed the attraction-repulsion system
$\left\{\begin{array}[]{rrll}&&{u}_{t}=\nabla\cdot(D({u})\nabla{u})-\chi\nabla\cdot({u}{\nabla{v}})+\xi\nabla\cdot({u}{\nabla{w}}),\\ &&\tau_{1}{v}_{t}=d_{1}\Delta{v}+\alpha{u}-\beta{v},\\ &&\tau_{2}{w}_{t}=d_{2}\Delta{w}+\gamma{u}-\delta{w}.\end{array}\right.$ (1.17)
For the case $\tau_{1}=1$ and $\tau_{2}=0$, Lin et. al. [29] proved the existence and global bounded solution with $D({u})=D_{0}{u}^{-\theta}$, under the condition that $\theta<\frac{2}{n}-1$. By using Lyapunov functional, the finite time blow up is also studied for the radially symmetric solutions with $\theta=\frac{2}{n}-1$ and $n=3$. In addition, blow-up of the solution to the sytem (1.17) may exist, if $\theta>\frac{2}{n}-1$ and $\chi\alpha-\xi\gamma>0$ with $n\geq 3$. On the contrary, if $\xi\gamma-\chi\alpha>0$ or $m>2-\frac{2}{n}$, Wang [40] discussed the global bounded solutions to the system (1.17) for the non-degenerate diffusion with $D({u})\geq C_{D}{u}^{m-1}$, $m\geq 1$. Additionally, the global bounded weak solution for the case of degenerate diffusion is proved under the same condition. If $\chi\alpha-\xi\gamma<0$, global bounded classical solution to the system (1.17) has been studied by Guo et. al. [13], while, if $\chi\alpha-\xi\gamma>0$, the solution of the system is bounded whenever $\|{u}_{0}\|_{{\mathcal{L}^{1}}(\Omega)}<\frac{4\pi}{\chi\alpha-\xi\gamma}$. On the other hand, if $\|{u}_{0}\|_{{\mathcal{L}^{1}}(\Omega)}>\frac{4\pi}{\chi\alpha-\xi\gamma}$, the solution may blow-up.
For the fully parabolic case, Jin et. al. [22] improved the conditions of Liu et. al. [31] in one space dimension. They also proved that, when the attractive and repulsive chemical signals have the same degradation with attraction dominating the repulsion, the solution of the system (1.17) converges as $t\to\infty$, algebraically. Conversely, Jin [20] showed the existence global bounded classical solution when the repulsion dominates the attraction in two space dimension. Further, they proved that the weak solution in three dimensional settings was discussed with large initial data. Based on the entropy-like inequality and coupled estimate techniques for global bounded classical solutions in two dimensional case, see Liu et. al. [30]. If $\xi\gamma=\chi\alpha$, the existence of global solutions to the system (1.17) as well as the asymptotic stability results were presented by Jin et. al. [21] for the case $n=2,3$. An interesting fact to note here that, this solution converge like the heat kernel as $t\to\infty$. Whenever $m>1$, where $D({u})\geq C({u}+\epsilon)^{m-1}$ with $\epsilon>0$, the global bounded classical solutions is examined by Wang et. al. [41]. They also proved that, the weak solution exists for the case $\epsilon=0$ under the same assumptions. If $\|{u}_{0}\|_{{\mathcal{L}^{1}}(\Omega)}<\frac{1}{C(\Omega)\chi\alpha}$, the existence of global bounded classical solutions and the asymptotic behavior to the system (1.17) is given by Lin et. al. [28] based on the entropy type inequality. In two dimensional case, Jin et. al. [23] proved the global existence of classical solutions under some conditions on the parameters and provide the asymptotic results with large initial data. In addition, if $\frac{\xi\gamma}{\chi\alpha}>\max\{\frac{\beta}{\delta},\frac{\delta}{\beta}\}$, the solution of the system (1.17) converge in terms of exponential. For more details, see [26, 43, 44, 46].
An attraction-repulsion chemotaxis system with logistic source is given by
$\left\{\begin{array}[]{rrll}&&{u}_{t}=\nabla\cdot(D({u})\nabla{u})-\chi\nabla\cdot({u}{\nabla{v}})+\xi\nabla\cdot({u}{\nabla{w}})+f({u}),\\ &&\tau{v}_{t}=d_{1}\Delta{v}+\alpha{u}-\beta{v},\\ &&\tau{w}_{t}=d_{2}\Delta{w}+\gamma{u}-\delta{w}.\end{array}\right.$ (1.21)
For the parabolic-elliptic case, where $D({u})\geq C_{D}({u}+\zeta)^{m-1}$, $m\geq 1$ and $f({u})\leq a-b{u}^{\eta}$, $\eta>1$ Wang [39] showed the existence of global bounded classical solutions to the system (1.21) for the case $\zeta>0$, and also, when either $\xi\gamma-\chi\alpha>0$ or logistic damping is large or diffusion is sufficiently large enough. Further, the system has weak solutions for the case $\zeta=0$ under the same assumptions. In addition, the asymptotic behavior has been studied to the system (1.21) with particular logistic source. Under some suitable conditions on the logistic source and the parameters, Zhang et. al. [48] obtained the global bounded classical solution as well as the weak solution to the system (1.21) with $D({u})\geq-C{u}(1+{u})$ and the large time behavior for $f({u})=\mu{u}(1-{u})$ respectively. Zhao et. al. [49] investigated the global bounded solution when $f({u})=r{u}-\mu{u}^{2}$ under the condition that $\chi\alpha-\xi\gamma>0$ and $\mu=\frac{1}{3}(\chi\alpha-\xi\gamma)$ with $D(u)\geq C_{D}u^{m-1}$, $m\geq 1$. Further, the exponentially convergence result is provided for the case when $D({u})=1$ and $n\geq 3$. On the other hand, they studied the boundedness and blow-up for the case when $D({u})=0$ and $n\geq 2$. When $D({u})\geq C_{D}({u}+1)^{m-1}$, $m\geq 1$ and $f({u})\leq a-b{u}^{\eta}$, $\eta>1$, Yan et. al. [45] proved the global bounded solution to the system (1.21) and this solution may blow-up whenever $\|{u}_{0}\|_{{\mathcal{L}^{1}}(\Omega)}>\frac{8\pi}{\chi\alpha-\xi\gamma}$ and $\chi\alpha-\xi\gamma>0$ under the specific functions in two dimensional settings. For the recent article see, [10].
Li et. al. [27] proved the system (1.21) has global bounded solutions for both the cases $\tau=0,1$ with $f({u})\leq a-b{u}^{\rho}$. For the case $\tau=0$ and $\xi\gamma-\chi\alpha>0$, there exits a bounded classical solution to the system (1.21) whenever $\rho\geq 1$. On the other hand, if $\xi\gamma-\chi\alpha<0$, still the result is valid whenever $\rho$ is strong enough. In one and two space dimensions, the author established the same results when $\tau>0$. For the fully parabolic case, Zheng et. al. [50] proved the existence of global bounded classical solution to the system (1.21) with $f({u})\leq a{u}-\mu{u}^{2}$ under some suitable conditions and this result improved the result of Li et. al. [27] provided $n\geq 3$ and $\beta=\delta$. In three dimensional settings, for $\beta,\delta\geq\frac{1}{2}$ and $\mu\geq\max\{(\frac{41}{2}\chi\alpha+9\xi\gamma)^{\eta},(9\chi\alpha+\frac{41}{2}\xi\gamma)^{\eta}\}$, the global bounded classical solutions to (1.21) with $f({u})={u}-\mu{u}^{\eta+1}$, $\eta>1$ has been studied by Li et. al. [25] and this solution converges to the steady state as $t\to\infty$. If, $\chi\alpha=\xi\gamma$, Shi et. al. [38] established the global bounded classical solutions to the system (1.21) with $f({u})\leq a-b{u}^{\theta}$ under $\theta$ and $n$ satisfying some suitable conditions. If $\mu>\frac{\chi^{2}\alpha^{2}(\beta-\delta)^{2}}{8\delta\beta^{2}}$, the exponential stability is discussed with $f({u})=\mu{u}(1-{u})$ for $n\leq 9$ by constructing appropriate Lyapunov functional. Suppose $D({u})\geq C{u}^{m-1}$ and also $m>\frac{6}{5}$, the global bounded classical solution is obtained by Zeng [47]. On the other hand, same results holds for $D({u})\leq C({u}^{m-1}+1)$ when $m\in(1,\frac{6}{5}]$. For more recent articles, one can refer [6, 14, 36].
Attraction-repulsion chemotaxis system with nonlinear signal production is studied with $f(u)\leq u(a-bu^{s})$, $f(0)\geq 0$ by Hong et. al. [16].
$\left\{\begin{array}[]{rrll}&&{u}_{t}=\Delta{u}-\chi\nabla\cdot({u}{\nabla{v}})+\xi\nabla\cdot({u}{\nabla{w}})+f({u}),\\ &&0=\Delta{v}+\alpha{u}^{k}-\beta{v},\\ &&0=\Delta{w}+\gamma{u}^{l}-\delta{w}.\end{array}\right.$ (1.25)
The existence of global solution is proved under some conditions on the attraction with max$\{l,s,\frac{2}{n}\}>k$. They also showed that the size of the coefficients will affect the boundedness of the solution and these results improved the results of [48]. Liu et. al. [32] established the finite time blowup results to the following system with nonlinear production
$\left\{\begin{array}[]{rrll}&&{u}_{t}=\Delta{u}-\chi\nabla\cdot({u}{\nabla{v}})+\xi\nabla\cdot({u}{\nabla{w}}),\\ &&0=\Delta{v}-\mu_{1}(t)+f_{1}(u),\\ &&0=\Delta{w}-\mu_{2}(t)+f_{2}(u),\end{array}\right.$ (1.29)
where, $f_{1}(s)=s^{\gamma_{1}}$, $f_{2}(s)=s^{\gamma_{2}}$, $\mu_{1}(t)=\frac{1}{|\Omega|}\int f_{1}(u)$ and $\mu_{2}(t)=\frac{1}{|\Omega|}\int f_{2}(u)$. The system (1.29) blowsup in finite time when $\gamma_{1}>\gamma_{2}$ and $\gamma_{1}>\frac{2}{n}$. In addition, global bounded solution is proved, for $\gamma_{1}<\frac{2}{n}$ with suitable initial data. The behaviour of the solution for the case $\gamma_{1}>\frac{2}{n}$ and $\gamma_{1}<\gamma_{2}$ is still an open question. Zhou et. al. [51] improved the condition for the global bounded solution to the system (1.25) which has already been proved by Hong et. al. [16] and also established the asymptotic behavior of the global solution to the system. The large time behavior of attraction repulsion system with nonlinear secretion was studied by Ren et. al. [37]. Here, the convergence rate of the solution is either exponential or algebraical.
For the study of nonlocal terms in the chemotaxis system, recently, Negreanu et al. [34] studied the following chemotaxis system with a nonlocal source
$\left\{\begin{array}[]{rrll}&&u_{t}=\Delta u-\chi\nabla\cdot(u^{m}\nabla v)+u(a_{0}-a_{1}u^{\alpha}+a_{2}\int_{\Omega}u^{\alpha}\mathrm{d}x),\\ &&v_{t}=\Delta v-v+u^{\gamma},\end{array}\right.$
If $\alpha\geq 1$, $m>1$, $\gamma\geq 1$, $\alpha+a>m+\gamma$, $a_{1}>0$, $a_{1}-a_{2}|\Omega|>0$ and under some suitable conditions on the initial data, the authors proved the global existence of solutions to the considered system. Moreover, the solution converges to the stead state ${u^{*}=\frac{a_{0}^{\frac{1}{\alpha}}}{(a_{1}-a_{2}|\Omega|)^{\frac{1}{\alpha}}},v^{*}=(u^{*})^{\gamma}}$.
Chiyo et al. [7] investigated the following system
$\left\{\begin{array}[]{rrll}&&u_{t}=\Delta u-\chi\nabla\cdot(u\nabla v)+au^{\alpha}-bu^{\alpha}\int_{\Omega}u^{\beta},\\ &&v_{t}=\Delta v-v+u,\end{array}\right.$
for $\tau=1$ and $\alpha,\beta\geq 1$. The authors established the uniform boundedness of classical solutions under two key regimes: the subquadratic case, $21\leq\alpha<$ 2 and $\beta>\frac{n+4}{2}-\alpha$; the superquadratic case, $\beta>\frac{n}{2},2\leq\alpha<1+\frac{2\beta}{n}$. This work extends the analysis of the fully parabolic case previously studied in [4].
Funtes et al. [12] studied the following system
$\left\{\begin{array}[]{rrll}&&u_{t}=\nabla\cdot((u+1)^{m_{1}-1}\nabla u-\chi u(u+1)^{m_{2}-1}\nabla v)+B(u,\nabla v),\\ &&v_{t}=\Delta v-v+f(u),\end{array}\right.$
where the term $B(u,\nabla u)$ takes one of the following forms $au^{\alpha}-bu^{\beta}-c\int_{\Omega}u^{\delta}$ or $au^{\alpha}-bu^{\alpha}\int_{\Omega}u^{\beta}-c|\nabla u|^{\delta}$. For each case, the authors derived sufficient conditions on the system parameters that ensure the existence of globally bounded classical solutions. This work extends previous analyses from [7, 4].
Ren [35] studied the existence of global bounded classical solutions to a class of attraction-repulsion chemotaxis systems with nonlocal terms. Hu [19] discussed the global existence of classical solutions to a two-species attraction-repulsion chemotaxis system with nonlocal terms and additionally studied the asymptotic stability under suitable Lyapunov functionals. In related work on nonlocal terms in chemotaxis systems, Columbu et al. [9] established global boundedness results for classical solutions of an attraction-repulsion chemotaxis model. Further developments can be found in [8].
To the best of our knowledge, no prior studies have investigated the attraction-repulsion chemotaxis system incorporating a nonlocal logistic source with sublinear production terms. Motivated by this gap, we introduce a nonlocal source term into the attraction-repulsion framework while accounting for sublinear production kinetics. Our analysis examines how these modifications influence the existence of solutions and their long-term dynamics.
Motivated by the above research works, main purpose of this paper is to establish the existence of the global classical solution to the system (1.6) with suitable conditions on $k$, $l$ and $m$, which is uniformly bounded in $\Omega\times(0,\infty)$ for $n\geq 2$. More to the point, this study is to examine the simultaneous effects of both attraction and repulsion mechanism in the presence of nonlocal source and sublinear production terms. Moreover, some specific comparisons are made for different values of parameters in numerical simulations.
Our article is organized as follows: Section 2, contains some basic inequalities, lemmas and the local existence of classical solution. Section 3 deals with boundedness and global existence of classical solution to the system (1.6). Section 4 is devoted to the presentation and discussion of our numerical examples. Finally, section 5 gives the conclusion of our work.
Theorem 1.1.
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ is an open, bounded domain with smooth boundary and $q>n$. Moreover, the function $f$ fulfills (1.10) with $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$. Assume that, there exists $p>\frac{n}{2}$, then for any initial data $({u}_{0},{v}_{0},{w}_{0})$ satisfy (1.9), then the system (1.6) possesses a unique classical solution $({u},{v},{w})$ which is uniformly bounded in the sense that
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+\big{\|}{v}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}+\big{\|}{w}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\leq C,\qquad\forall\,t>0,$
where the constant $C$ is positive.
Remark 1.1.
For the case $n=1$, the solution (u, v, w) exists globally by Lemma 2.4 with $p=1$, without any additional restrictions on the parameters.
2 Preliminaries and local existence
We recall some useful inequalities and key lemma which we are going to use in the sequel. In this section, we begin with the local existence of solutions to system (1.6) which is standard and its proof is based on the ideas of [18].
Lemma 2.1 (Local Existence).
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 2,$ is an open, bounded domain with smooth boundary and assume that $q>n$ and $M_{0}>0$. Moreover, the function $f$ fulfills (1.10) and $k,m\geq 1$, then for each non-negative initial data satisfies (1.9), there exists $T_{\mathrm{max}}\in(0,\infty]$ such that a uniquely determined triple $({u},{v},{w})$ of nonnegative functions
${u}$ $\in{\mathcal{C}^{0}}\left(\overline{\Omega}\times\left.\left[0,T_{\mathrm{max}}\right.\right)\right)\cap{\mathcal{C}^{2,1}}\left(\overline{\Omega}\times\left(0,T_{\mathrm{max}}\right)\right),$
${v},{w}$ $\in{\mathcal{C}^{0}}\left(\overline{\Omega}\times\left.\left[0,T_{\mathrm{max}}\right.\right)\right)\cap{\mathcal{C}^{2,1}}\left(\overline{\Omega}\times\left(0,T_{\mathrm{max}}\right)\right)\cap{\mathcal{L}^{\infty}_{loc}}\left(\left.\left[0,T_{\mathrm{max}}\right.\right);{\mathcal{W}^{1,q}}(\Omega)\right),$
solving (1.6) classically in $\Omega\times(0,T_{\mathrm{max}})$. Furthermore, if $T_{\mathrm{max}}<\infty$, then
$\lim_{t\to T_{\mathrm{max}}}\Big{(}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+\big{\|}{v}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}+\big{\|}{w}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\Big{)}=\infty.$ (2.1)
Proof.
The proof can be derived by the standard arguments involving the parabolic regularity theory. For the proof we refer to [18] and [11]. In addition, the non-negativity of the solution in $\Omega\times(0,T_{\mathrm{max}})$ is comes from the maximum principle along with (1.9). ∎
Lemma 2.2.
(See [3]) For $\int_{\Omega}{u}(\cdot,0)=M_{0}>0$, the mass $\int_{\Omega}{u}(x,t)\mathrm{d}x=M(t)$ satisfies
$\min\{1,M_{0}\}\leq M(t)\leq\max\{1,M_{0}\}.$
Furthermore, we have the following decay estimates
$|1-M(t)|\leq|1-M_{0}|e^{-\min\big{\{}1,M_{0}^{m}\big{\}}t}.$
Lemma 2.3 (Maximal Sobolev regularity [5, 15]).
Let $r\in(1,\infty)$ and $T\in(0,\infty)$. Consider the following evolution equation
$\left\{\begin{array}[]{rrll}&&y_{t}=\Delta y-y+g,&x\in\Omega,\>t>0,\\ &&\frac{\partial y}{\partial\nu}=0,&x\in\partial\Omega,\>t>0,\\ &&y(x,0)=y_{0}(x),&x\in\Omega.\end{array}\right.$
For each $y_{0}\in{\mathcal{W}^{2,r}}(\Omega)$ such that $\frac{\partial y(\cdot,s_{0})}{\partial\nu}=0$ on $\partial\Omega$ and any $g\in{\mathcal{L}^{r}}\big{(}(0,T);{\mathcal{L}^{r}}(\Omega)\big{)}$, there exists a unique solution
$y\in{\mathcal{W}^{1,r}}\big{(}(0,T);{\mathcal{L}^{r}}(\Omega)\big{)}\cap{\mathcal{L}^{r}}\big{(}(0,T);{\mathcal{W}^{2,r}}(\Omega)\big{)}.$
Moreover, there exists $C_{r}>0$ such that
$\int_{s_{0}}^{T}\int_{\Omega}y^{r}+\int_{s_{0}}^{T}\int_{\Omega}|y_{t}|^{r}+\int_{s_{0}}^{T}\int_{\Omega}|\Delta y|^{r}\leq C_{r}\int_{s_{0}}^{T}\int_{\Omega}|g|^{r}+C_{r}\int_{\Omega}y_{0}^{r}+C_{r}\int_{\Omega}|\Delta y_{0}|^{r}.$
If $s_{0}\in(0,T)$ and $y(\cdot,s_{0})\in{\mathcal{W}^{2,r}}(\Omega)$ with $\frac{\partial y(\cdot,s_{0})}{\partial\nu}=0$ on $\partial\Omega$ we have
$\int_{s_{0}}^{T}\int_{\Omega}e^{sr}|\Delta y|^{r}\leq C_{r}\int_{s_{0}}^{T}\int_{\Omega}e^{sr}|g|^{r}+C_{r}\int_{\Omega}y(\cdot,s_{0})^{r}+C_{r}\int_{\Omega}|\Delta y(\cdot,s_{0})|^{r}.$
Lemma 2.4 (Extensibility criterion).
Let $n\geq 1$, $M_{0}>0$ and the initial data ${u}_{0}$, ${v}_{0}$ and ${w}_{0}$ satisfy (1.9) for some $q>n$. Moreover, the function g fulfill (1.10) with $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$. Suppose that there exists $p\geq 1$ such that $p>\frac{n}{2}$ and
$\sup\limits_{t\in(s_{0},T_{\mathrm{max}})}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{p}}(\Omega)}<\infty,$
then, we have
$\sup\limits_{t\in(s_{0},T_{\mathrm{max}})}\Big{(}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+\big{\|}{v}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}+\big{\|}{w}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\Big{)}<\infty.$ (2.2)
Proof.
Let $q>n$ and for each fixed $p>\frac{n}{2}$ there holds
$\frac{np}{(n-p)_{+}}=\left\{\begin{array}[]{rrll}&&\infty,&\quad\mbox{if}\quad p\geq n,\\ \\ &&\frac{np}{n-p}>n,&\quad\mbox{if}\quad\frac{n}{2}<p<n,\end{array}\right.$ (2.6)
[…]
${v}(\cdot,t)=e^{-\alpha t}e^{d_{2}t\Delta}{v}_{0}+\int^{t}_{0}e^{-\alpha(t-s)}e^{d_{2}(t-s)\Delta}f\big{(}u(\cdot,t)\big{)}\mathrm{d}s.$
Now,
$\big{\|}\nabla{v}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}\leq$ $\>e^{-\alpha t}\big{\|}\nabla e^{d_{2}t\Delta}{v}_{0}\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
$+K\int^{t}_{0}e^{-\alpha(t-s)}\Big{\|}\nabla e^{d_{2}(t-s)\Delta}{u}^{l}(\cdot,s)\Big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}\mathrm{d}s.$
By use of the Neumann heat semigroup [42] with $q_{0}\sigma\leq q$, we attain
$\big{\|}\nabla{v}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}\leq$ $\>C_{1}e^{-\alpha t}\|{v}_{0}\|_{{\mathcal{W}^{1,q}}(\Omega)}$
$+C_{2}\int^{t}_{0}e^{-\alpha(t-s)}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{p}-\frac{1}{q_{0}\sigma}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{l}(\cdot,s)\Big{\|}_{{\mathcal{L}^{p}}(\Omega)}\mathrm{d}s,$ (2.7)
where $C_{1}$ and $C_{2}$ are positive constants. Here, we can guarantee that $\frac{1}{2}+\frac{n}{2}\left(\frac{1}{p}-\frac{1}{q_{0}\sigma}\right)<1$, because of our assumption $q_{0}\sigma<\frac{np}{(n-p)}$ and $q_{0}\sigma\leq q$. Now, adopting the Gamma function, we ensure $\int_{0}^{\infty}e^{-\alpha\psi}\left(1+\psi^{-\frac{1}{2}-\frac{n}{2}(\frac{1}{p}-\frac{1}{q_{0}\sigma})}\right)e^{-\lambda\psi}<\infty$. Using the Interpolation inequality with Lemma 3.2, to get
$\big{\|}{u}^{l}(\cdot,t)\big{\|}_{{\mathcal{L}^{p}}(\Omega)}\leq|\Omega|^{\frac{1-l}{p}}\left(\int_{\Omega}u(\cdot,t)^{p}\right)^{\frac{l}{p}}\leq C_{3}\|{u}(\cdot,t)\big{\|}^{l}_{{\mathcal{L}^{p}}(\Omega)},$ (2.8)
where $l\in(0,1)$. Substituting (2.8) in to (2.7), we arrive at
$\big{\|}\nabla{v}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{1}\|{v}_{0}\|_{{\mathcal{W}^{1,q}}(\Omega)}+C_{4},$
where $C_{4}$ is positive constant. Finally, we obtain
$\big{\|}\nabla{v}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{5},\qquad\qquad\forall\,t\in(s_{0},T_{\mathrm{max}}).$ (2.9)
Like wise we obtain,
$\big{\|}{\nabla{w}}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{6},\qquad\qquad\forall\,t\in(s_{0},T_{\mathrm{max}}).$ (2.10)
where $C_{5}$ and $C_{6}>0$. We can attain the following representation by adopting the variation of constants formula
${u}(\cdot,t)=$ $e^{d_{1}t\Delta}{u}(\cdot,0)-\chi\int^{t}_{0}e^{d_{1}(t-s)\Delta}\nabla\cdot\Big{(}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\Big{)}\mathrm{d}s$
$+\xi\int^{t}_{0}e^{d_{1}(t-s)\Delta}\nabla\cdot\Big{(}{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\Big{)}\mathrm{d}s$
$+\mu\int^{t}_{0}e^{d_{1}(t-s)\Delta}{u}^{m}(\cdot,s)\Big{(}1-\int_{\Omega}{u}(\cdot,s)\mathrm{d}x\Big{)}\mathrm{d}s.$ (2.11)
The above representation yields
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq$ $\>\Big{\|}e^{d_{1}t\Delta}{u}(\cdot,0)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+\chi\int^{t}_{0}\Big{\|}e^{d_{1}(t-s)\Delta}\nabla\cdot\Big{(}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\Big{)}\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s$
$+\xi\int^{t}_{0}\Big{\|}e^{d_{1}(t-s)\Delta}\nabla\cdot\Big{(}{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\Big{)}\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s$
$+\mu\int^{t}_{0}\left\|e^{d_{1}(t-s)\Delta}{u}^{m}(\cdot,s)\left(1-M(t)\right)\right\|_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s,$
for all $t\in(0,T_{\mathrm{max}})$. Now, applying the Neumann heat semigroup [42] with $C_{7}$ and $C_{8}>0$ satisfying
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq$ $\>\big{\|}{u}(\cdot,0)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}$
$+C_{7}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\Big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s$
$+C_{8}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\Big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s$
$+|1-M_{0}|C_{9}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{m}(\cdot,s)\Big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s,$ (2.12)
where $\frac{1}{2}+\frac{n}{2q_{0}}<1$ because of $q_{0}>n$. Employing the Gamma function again, the integral $\int_{0}^{\infty}\left(1+\psi^{-\frac{1}{2}-\frac{n}{2q_{0}}}\right)e^{-\lambda\psi}<\infty.$ The Hölder’s inequality and the Interpolation inequality gives us that
$\big{\|}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}$ $\leq\big{\|}{u}^{k}(\cdot,s)\big{\|}_{L^{q_{0}\widehat{\sigma}}(\Omega)}\;\big{\|}{\nabla{v}(\cdot,s)}\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
$\leq\big{\|}{u}(\cdot,s)\big{\|}^{k}_{L^{q_{0}\widehat{\sigma}k}(\Omega)}\;\big{\|}{\nabla{v}(\cdot,s)}\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
$\leq\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{1}}(\Omega)}^{k\zeta_{2}}\;\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{k(1-\zeta_{2})}\;\big{\|}{\nabla{v}(\cdot,s)}\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
$\leq\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{1}}(\Omega)}^{\frac{1}{q_{0}\widehat{\sigma}}}\;\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{k-\frac{1}{q_{0}\widehat{\sigma}}}\;\big{\|}{\nabla{v}(\cdot,s)}\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
$\leq C_{10}\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{k-\frac{1}{q_{0}\widehat{\sigma}}}.$
Now, using the Young’s inequality with $q_{0}\widehat{\sigma}>\frac{n}{2}$, gives
$\big{\|}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\big{\|}_{L^{q_{0}}(\Omega)}$ $\leq\frac{1}{6}\big{\|}{u}(\cdot,s)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{11}.$ (2.13)
Similarly, we can attain
$\big{\|}{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\big{\|}_{L^{q_{0}}(\Omega)}$ $\leq\frac{1}{6}\big{\|}{u}(\cdot,s)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{12},$ (2.14)
where $\widehat{\sigma}$ is the dual exponent of $\sigma$ and ${\zeta_{2}=\frac{1}{q_{0}\widehat{\sigma}k}\in(0,1)}$, for all $s\in(s_{0},T_{\mathrm{max}})$. Also,
$\big{\|}{u}^{m}(\cdot,s)\big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}=$ $\Big{\|}{u}(\cdot,s)\Big{\|}^{m}_{L^{q_{0}m}(\Omega)}$
$\leq$ $\Big{\|}{u}(\cdot,s)\Big{\|}^{m\zeta_{3}}_{{\mathcal{L}^{1}}(\Omega)}\>\Big{\|}{u}(\cdot,s)\Big{\|}^{m(1-\zeta_{3})}_{{\mathcal{L}^{\infty}}(\Omega)}$
$\leq$ $C_{13}\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{m-\frac{1}{q_{0}}},$
where ${\zeta_{3}=\frac{1}{q_{0}m}\in(0,1)}$, for all $s\in(s_{0},T_{\mathrm{max}})$ and $C_{13}>0$. Once again the Young’s inequality with $q_{0}>\frac{n}{2}$, gives
$\big{\|}{u}^{m}(\cdot,s)\big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\leq$ $C_{13}\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{m-\frac{1}{q_{0}}}\leq\frac{1}{6}\big{\|}{u}(\cdot,s)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{14}.$ (2.15)
Inserting (2.13)-(2.15) in (2.12), it follows that
$\sup_{t\in(s_{0},T)}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq\frac{1}{2}\sup_{t\in(s_{0},T)}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{15},$
for all $T\in(s_{0},T_{\mathrm{max}})$. Finally, we conclude that
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq C_{16},\qquad\qquad\qquad\forall\,t\in(s_{0},T_{\mathrm{max}}).$ (2.16)
where the constant $C_{16}$ is positive. ∎
3 Global existence of classical solutions
In this section, we prove the global existence and boundedness of the solution to (1.6). First we derive ${\mathcal{L}^{p}}(\Omega)$ norm for ${u}$. Given any $s_{0}\in(0,T_{\mathrm{max}})$ such that $s_{0}<1$, Lemma 2.1 gives ${u}(\cdot,s_{0}),{v}(\cdot,s_{0}),{w}(\cdot,s_{0})\in{\mathcal{C}^{2}(\overline{\Omega})}$ with $\frac{\partial{v}(\cdot,s_{0})}{\partial\nu}=0$ and $\frac{\partial{w}(\cdot,s_{0})}{\partial\nu}=0$, we pick $C>0$ such that
$\left\{\begin{array}[]{rrll}&&\sup\limits_{0\leq s\leq s_{0}}\|{u}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\quad\sup\limits_{0\leq s\leq s_{0}}\|{v}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\quad\sup\limits_{0\leq s\leq s_{0}}\|{w}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\\ \\ &&\|\Delta{v}(\cdot,s_{0})\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C\quad\mathrm{and}\quad\|\Delta{w}(\cdot,s_{0})\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C.\end{array}\right.$ (3.4)
Next, we derive boundedness in $t\in(s_{0},T_{\mathrm{max}})$.
Lemma 3.1.
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 1$ is an open, bounded domain with smooth boundary. Assume that $p+m-1>p^{\prime}>1$ and $M_{0}<1$. If $m$ take the values,
11 11 $<m<1+\frac{2}{n},\hskip 56.9055pt\mathrm{if}\quad n\geq 3,$
11 11 $<m<2,\hskip 79.6678pt\mathrm{if}\quad n=1,2,$
then we have
$\int_{\Omega}(1-M_{0}){{u}}^{p+m-1}$ $\leq\frac{2d_{1}(p-1)}{p^{2}\mu}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}+C,$ (3.5)
where $C>0$.
Proof.
Using the Hölder’s inequality
$\int_{\Omega}{{u}}^{p+m-1}=$ $\int_{\Omega}{u}^{\frac{p}{2}\frac{2\gamma(p+m-1)}{p}}{u}^{\frac{p}{2}\frac{2(1-\gamma)(p+m-1)}{p}}$
$\leq$ $\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2\gamma(p+m-1)}{p}\frac{rp}{2\gamma(p+m-1)}}\right)^{\frac{2\gamma(p+m-1)}{rp}}$
×$\times\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2(1-\gamma)(p+m-1)}{p}\frac{p}{2(1-\gamma)(p+m-1)}\frac{2p^{\prime}}{p}}\right)^{\frac{2(1-\gamma)(p+m-1)}{p}\frac{p}{2p^{\prime}}}$
$\leq$ $\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{r}\right)^{\frac{1}{r}\frac{2\gamma(p+m-1)}{p}}\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2p^{\prime}}{p}}\right)^{\frac{p}{2p^{\prime}}\frac{2(1-\gamma)(p+m-1)}{p}},$
from the above, we obtain the estimate
$\int_{\Omega}{{u}}^{p+m-1}\leq$ $\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{r}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}.$ (3.6)
Using the Sobolev embedding theorem, we get the estimate from (3.6)
$\int_{\Omega}{{u}}^{p+m-1}\leq$ $C_{1}\Bigg{(}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\gamma}_{{\mathcal{L}^{2}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{(1-\gamma)}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}+\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{\gamma}_{{\mathcal{L}^{2}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{(1-\gamma)}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}\Bigg{)}^{\frac{2(p+m-1)}{p}}$
$\leq$ $C_{1}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{2}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}+C_{1}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{2}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}.$ (3.7)
Where $\gamma$ is the exponent from the Hölder’s inequality, such that
$\frac{2\gamma(p+m-1)}{rp}+\frac{(1-\gamma)(p+m-1)}{p^{\prime}}$ $=1$
$\frac{2\gamma}{rp}+\frac{(1-\gamma)}{p^{\prime}}$ $=\frac{1}{p+m-1}$
$\frac{\gamma}{r}+\frac{p}{2p^{\prime}}-\frac{\gamma p}{2p^{\prime}}$ $=\frac{p}{2(p+m-1)}$
$\gamma\left[\frac{1}{r}-\frac{p}{2p^{\prime}}\right]$ $=\frac{p}{2(p+m-1)}-\frac{p}{2p^{\prime}},$
we have,
$\gamma$ $=\frac{\frac{p}{2p^{\prime}}-\frac{p}{2(p+m-1)}}{\frac{p}{2p^{\prime}}-\frac{1}{r}}\in(0,1),$ (3.8)
and $r$ satisfies
$r=\left\{\begin{array}[]{rrll}&&\frac{2n}{n-2},&\quad\mbox{if}\quad n\geq 3,\\ \\ &&\left(\frac{2(p+m-1)}{p},\infty\right),&\quad\mbox{if}\quad n=2,\\ \\ &&\infty,&\quad\mbox{if}\quad n=1.\end{array}\right.$
Using the Young’s inequality to the second term in R.H.S of (3.7), we obtain
$C_{1}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{2}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}\leq$ $\>\frac{2d_{1}(p-1)}{p^{2}\mu}\Bigg{(}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{2}}(\Omega)}\Bigg{)}^{\frac{p}{\gamma(p+m-1)}}$
$+C(d_{1},p,\mu)\Bigg{(}C_{1}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}\Bigg{)}^{\frac{1}{1-\frac{\gamma(p+m-1)}{p}}}$
$\leq$ $\>\frac{2d_{1}(p-1)}{p^{2}\mu}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{2}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{a}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)},$ (3.9)
where $a=\frac{2(1-\gamma)(p+m-1)}{p}\frac{1}{1-\frac{\gamma(p+m-1)}{p}}$. Take $p^{\prime}>\frac{n(m-1)}{2}$ with simple computation, we ensure that
$\frac{2\gamma(p+m-1)}{p}<$ $\frac{\frac{pn}{p^{\prime}}+2-n}{\frac{1}{2}\left(\frac{pn}{p^{\prime}}+2-n\right)}<2.$
Substituting (3.9) in (3.7), we obtain
$\int_{\Omega}{{u}}^{p+m-1}\leq$ $\frac{2d_{1}(p-1)}{p^{2}\mu}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{2}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{a}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}+C_{1}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{2}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}.$ (3.10)
In addition, we have the following estimates
$(i)\quad$ $\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{a}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}=\Bigg{(}\int_{\Omega}\Big{(}{u}^{\frac{p}{2}}\Big{)}^{\frac{2p^{\prime}}{p}}\Bigg{)}^{\frac{p}{2p^{\prime}}a}=\Big{\|}{u}\Big{\|}^{\frac{pa}{2}}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)},$ (3.11)
$(ii)\quad$ $\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{2}}(\Omega)}=\Bigg{(}\int_{\Omega}\Big{(}{u}^{\frac{p}{2}}\Big{)}^{2}\Bigg{)}^{\frac{1}{2}\frac{2\gamma(p+m-1)}{p}}=\Big{\|}{u}\Big{\|}^{\gamma(p+m-1)}_{{\mathcal{L}^{p}}(\Omega)},$ (3.12)
$(iii)\quad$ $\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}=\Bigg{(}\int_{\Omega}\Big{(}{u}^{\frac{p}{2}}\Big{)}^{\frac{2p^{\prime}}{p}}\Bigg{)}^{\frac{p}{2p^{\prime}}\frac{2(1-\gamma)(p+m-1)}{p}}=\Big{\|}{u}\Big{\|}^{(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}.$ (3.13)
Substituting (3.11)-(3.13) into (3.10), we get
$\int_{\Omega}{{u}}^{p+m-1}\leq\frac{2d_{1}(p-1)}{p^{2}\mu}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{2}\Big{\|}{u}\Big{\|}^{\frac{pa}{2}}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}+C_{1}\Big{\|}{u}\Big{\|}^{\gamma(p+m-1)}_{{\mathcal{L}^{p}}(\Omega)}\>\Big{\|}{u}\Big{\|}^{(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}.$ (3.14)
On the other hand, using the Hölder’s inequality, we have
$\Big{\|}{u}\Big{\|}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq$ $C_{3}\Bigg{(}\int_{\Omega}{u}^{\theta p^{\prime}\frac{p+m-1}{\theta p^{\prime}}}\Bigg{)}^{\frac{\theta p^{\prime}}{p^{\prime}(p+m-1)}}\Bigg{(}\int_{\Omega}{u}^{(1-\theta)p^{\prime}\frac{1}{(1-\theta)p^{\prime}}}\Bigg{)}^{\frac{(1-\theta)p^{\prime}}{p^{\prime}}}$
$\leq$ $C_{3}\Big{\|}{u}\Big{\|}^{\theta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{(1-\theta)}_{{\mathcal{L}^{1}}(\Omega)},$
where $\theta=\frac{(p^{\prime}-1)(p+m-1)}{p^{\prime}(p+m-2)}\in(0,1)$. Similar way, we pick up
$\Big{\|}{u}\Big{\|}_{{\mathcal{L}^{p}}(\Omega)}\leq$ $C_{4}\Bigg{(}\int_{\Omega}{u}^{\eta p\frac{p+m-1}{\eta p}}\Bigg{)}^{\frac{\eta p}{p(p+m-1)}}\Bigg{(}\int_{\Omega}{u}^{(1-\eta)p\frac{1}{(1-\eta)p}}\Bigg{)}^{\frac{(1-\eta)p}{p}}$
$\leq$ $C_{4}\Big{\|}{u}\Big{\|}^{\eta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{(1-\eta)}_{{\mathcal{L}^{1}}(\Omega)},$
with $\eta=\frac{(p-1)(p+m-1)}{p(p+m-2)}\in(0,1)$. Hence, we obtain the following estimates from the above two inequalities,
$(i)\quad C_{2}\Big{\|}{u}\Big{\|}^{\frac{pa}{2}}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq C_{2}\Bigg{(}C_{3}\Big{\|}{u}\Big{\|}^{\theta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{1-\theta}_{{\mathcal{L}^{1}}(\Omega)}\Bigg{)}^{\frac{pa}{2}}\leq C_{5}\Big{\|}{u}\Big{\|}^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)},$ (3.15)
$(ii)\quad C_{1}\Big{\|}{u}\Big{\|}^{\gamma(p+m-1)}_{{\mathcal{L}^{p}}(\Omega)}\leq C_{1}\Bigg{(}C_{4}\Big{\|}{u}\Big{\|}^{\eta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{1-\eta}_{{\mathcal{L}^{1}}(\Omega)}\Bigg{)}^{\gamma(p+m-1)}\leq C_{6}\Big{\|}{u}\Big{\|}^{\eta\gamma(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)},$ (3.16)
$(iii)\quad\Big{\|}{u}\Big{\|}^{(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq\Bigg{(}C_{3}\Big{\|}{u}\Big{\|}^{\theta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{1-\theta}_{{\mathcal{L}^{1}}(\Omega)}\Bigg{)}^{(1-\gamma)(p+m-1)}\leq C_{7}\Big{\|}{u}\Big{\|}^{\theta(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}.$ (3.17)
Substituting (3.15)-(3.17) in (3.14), finally we achieve
$\int_{\Omega}{{u}}^{p+m-1}\leq$ $\frac{2d_{1}(p-1)}{p^{2}\mu}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{5}\Big{\|}{u}\Big{\|}^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C_{8}\Big{\|}{u}\Big{\|}^{\eta\gamma(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{\theta(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}$
$\leq$ $\frac{2d_{1}(p-1)}{p^{2}\mu}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{5}\Big{\|}{u}\Big{\|}^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C_{8}\Big{\|}{u}\Big{\|}^{[\eta\gamma+\theta(1-\gamma)](p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}.$ (3.18)
Again, using the Young’s inequality, the second term can be written as
$C_{5}\Big{\|}{u}\Big{\|}^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)}\leq\frac{M_{0}}{2}\Big{\|}{u}\Big{\|}^{\frac{pa\theta}{2}\frac{2(p+m-1)}{pa\theta}}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C(M_{0})C_{9}\leq\frac{M_{0}}{2}\Big{\|}{u}\Big{\|}^{p+m-1}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C_{10}.$ (3.19)
By simple calculations, we guarantee that $\frac{pa\theta}{2(p+m-1)}<1$. That is,
$\frac{(1-\gamma)(p+m-1)}{1-\frac{\gamma(p+m-1)}{p}}\frac{1-\frac{1}{p^{\prime}}}{1-\frac{1}{p+m-1}}\frac{1}{p+m-1}<$ 11 11
$(1-\gamma)(p+m-1)\left(1-\frac{1}{p^{\prime}}\right)<$ $\left(1-\frac{\gamma(p+m-1)}{p}\right)(p+m-2)$
$\frac{1}{p+m-1}+\frac{\gamma-1}{p^{\prime}}<$ $\gamma\left(1-\frac{p+m-1}{p}\right)+\frac{\gamma}{p}$
$\frac{p}{\gamma(p+m-1)}-\frac{p}{\gamma p^{\prime}}<$ $(2-m)-\frac{p}{p^{\prime}}$
$\gamma<$ $\frac{\frac{p}{p^{\prime}}-\frac{p}{p+m-1}}{\frac{p}{p^{\prime}}-(2-m)}.$ (3.20)
Once again, using the Young’s inequality to the third term of (3.18), the estimates gives
$C_{8}\Big{\|}{u}\Big{\|}^{[\eta\gamma+\theta(1-\gamma)](p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}\leq$ $\frac{M_{0}}{2}\left(\Big{\|}{u}\Big{\|}^{[\eta\gamma+\theta(1-\gamma)](p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}\right)^{\frac{1}{\eta\gamma+\theta(1-\gamma)}}+C(M_{0})C_{11}$
$\leq$ $\frac{M_{0}}{2}\Big{\|}{u}\Big{\|}^{p+m-1}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C_{12},$ (3.21)
where $\eta\gamma+\theta(1-\gamma)<1$ because of the definition of $\gamma$, $\theta$ and $\eta$. Substituting (3.19) and (3.21) in (3.18), we obtain that
$\int_{\Omega}{{u}}^{p+m-1}$ $\leq\frac{2d_{1}(p-1)}{p^{2}\mu}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{2}_{{\mathcal{L}^{2}}(\Omega)}+\frac{M_{0}}{2}\Big{\|}{u}\Big{\|}^{p+m-1}_{{\mathcal{L}^{p+m-1}}(\Omega)}+\frac{M_{0}}{2}\Big{\|}{u}\Big{\|}^{p+m-1}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C_{13},$
where the constant $C_{13}>0$. Finally, we obtain
$\int_{\Omega}(1-M_{0}){{u}}^{p+m-1}$ $\leq\frac{2d_{1}(p-1)}{p^{2}\mu}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}+C_{13}.$
Now, we will derive the condition to the first case. For $n\geq 3$ and $r=\frac{2n}{n-2}$, we have from (3.8)
$\gamma$ $=\frac{\frac{pn}{2p^{\prime}}-\frac{pn}{2(p+m-1)}}{\frac{pn}{2p^{\prime}}+1-\frac{n}{2}}=\frac{\frac{p}{p^{\prime}}-\frac{p}{p+m-1}}{\frac{p}{p^{\prime}}-1+\frac{2}{n}}.$ (3.22)
Comparing the gamma values (3.20) with (3.22), we acquire
$m<1+\frac{2}{n}.$
Next, for the case $n=2$, we have
$\gamma=\frac{\frac{p}{p^{\prime}}-\frac{p}{p+m-1}}{\frac{p}{p^{\prime}}-\frac{2}{r}}.$
Now (3.19) holds true for the case $n=2$ whenever $m<2-\frac{2}{r}<2$ and also if, $m<2$, (3.19) remains true for the case $n=1$. This completes the proof. ∎
Lemma 3.2.
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 1$ is an open, bounded domain with smooth boundary. Moreover, let $f$ fulfills (1.10) and for any $p>1$ and $M_{0}>0$, if $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$, then we have
$\hskip 85.35826pt\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{p}}(\Omega)}\leq C,\hskip 85.35826pt\forall\,t\in(s_{0},T_{\mathrm{max}}),$ (3.23)
for some $C>0$.
Proof.
From Lemma 2.2, $M_{0}\leq M(t)\leq 1$ for $M_{0}<1$. Multiply with $p{{u}}^{p-1}$, $p>1$, in the first equation of (1.6) and integrate over $\Omega$, we get
$p\int_{\Omega}{u}_{t}{{u}}^{p-1}=$ $d_{1}p\int_{\Omega}{{u}}^{p-1}\Delta{u}-p\chi\int_{\Omega}{{u}}^{p-1}\nabla\cdot\left({u}^{k}{\nabla{v}}\right)+p\xi\int_{\Omega}{{u}}^{p-1}\nabla\cdot\left({u}^{k}{\nabla{w}}\right)$
$+\mu p\int_{\Omega}{{u}}^{p-1}{u}^{m}\left(1-\int_{\Omega}{u}(x,t)\mathrm{d}x\right).$
The integration by parts, gives us that
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}\leq$ $-d_{1}p(p-1)\int_{\Omega}{{u}}^{p-2}|{\nabla{u}}|^{2}+\chi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{v}}$
$-\xi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{w}}+\mu p\int_{\Omega}{{u}}^{p+m-1}-\mu pM_{0}\int_{\Omega}{{u}}^{p+m-1}.$ (3.24)
Again use of integration by parts to the second and third term in R.H.S and using Lemma 3.1, leads to
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}\leq$ $\frac{-4d_{1}(p-1)}{p}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}-\chi p\frac{p-1}{p+k-1}\int_{\Omega}{{u}}^{p+k-1}{\Delta{v}}+\xi p\frac{p-1}{p+k-1}\int_{\Omega}{{u}}^{p+k-1}{\Delta{w}}$
$+\frac{2d_{1}(p-1)}{p}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}+C_{0}$
$\leq$ $\frac{-2d_{1}(p-1)}{p}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}+I_{1}+I_{2}+C_{0}.$ (3.25)
Let $\delta=\frac{np+2}{(p+k-1)n}>1,$ now applying the Young’s inequality with $\epsilon>0$, we get
$I_{1}=\left|-\chi p\frac{p-1}{p+k-1}\int_{\Omega}{{u}}^{p+k-1}{\Delta{v}}\right|\leq$ $\epsilon\int_{\Omega}{u}^{(p+k-1)\delta}+C_{1}\int_{\Omega}{|\Delta{v}|}^{\frac{\delta}{\delta-1}}$
$\leq$ $\epsilon\int_{\Omega}{u}^{\frac{np+2}{n}}+C_{1}\int_{\Omega}{|\Delta{v}|}^{\frac{np+2}{2-(k-1)n}},$ (3.26)
where $C_{1}=\left(\epsilon\delta\right)^{\frac{-1}{\delta-1}}\frac{\delta-1}{\delta}\left(\chi p\frac{p-1}{p+k-1}\right)^{\frac{\delta}{\delta-1}}$. Similarly for $I_{2}$
$I_{2}=\xi p\frac{p-1}{p+k-1}\int_{\Omega}{{u}}^{p+k-1}{\Delta{w}}\leq$ $\epsilon\int_{\Omega}{u}^{(p+k-1)\delta}+C_{2}\int_{\Omega}{|\Delta{w}|}^{\frac{\delta}{\delta-1}}$
$\leq$ $\epsilon\int_{\Omega}{u}^{\frac{np+2}{n}}+C_{2}\int_{\Omega}{|\Delta{w}|}^{\frac{np+2}{2-(k-1)n}},$ (3.27)
where $C_{2}=\left(\epsilon\delta\right)^{\frac{-1}{\delta-1}}\frac{\delta-1}{\delta}\left(\xi p\frac{p-1}{p+k-1}\right)^{\frac{\delta}{\delta-1}}$. Substituting. (3.26) and (3.27) in (3.25), we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}+\frac{d_{1}(p-1)}{p}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}\leq$ $\frac{-d_{1}(p-1)}{p}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}+2\epsilon\int_{\Omega}{u}^{\frac{np+2}{n}}$
$+C_{1}\int_{\Omega}{|\Delta{v}|}^{\frac{np+2}{2-(k-1)n}}+C_{2}\int_{\Omega}{|\Delta{w}|}^{\frac{np+2}{2-(k-1)n}}+C_{0}.$ (3.28)
Using the Gagliardo-Nirenberg inequity and the Young’s inequality, we achieve
$\int_{\Omega}{{u}}^{p}$ $=\left\|{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}\leq C_{3}\left(\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2b}_{{\mathcal{L}^{2}}(\Omega)}\,\,\left\|{u}^{\frac{p}{2}}\right\|^{2(1-b)}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}+\left\|{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}\right)$
$\leq\epsilon\left(\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2b}_{{\mathcal{L}^{2}}(\Omega)}\right)^{\frac{1}{b}}+C_{3}\,C(\epsilon)\left(\left\|{u}^{\frac{p}{2}}\right\|^{2(1-b)}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}\right)^{\frac{1}{1-b}}+C_{3}\left\|{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}$
$\leq\epsilon\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{4}\left\|{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}$
$\leq\epsilon\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{4}\Big{\|}{u}\Big{\|}^{p}_{{\mathcal{L}^{1}}(\Omega)}$
$\leq\frac{2-(k-1)n}{np+2}\frac{d_{1}(p-1)}{p}\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{5}$
$\leq\frac{2-(k-1)n}{np+2}\frac{d_{1}(p-1)}{p}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}+C_{5},$
where $b=\frac{\frac{p}{2}-\frac{1}{2}}{\frac{p}{2}+\frac{1}{n}-\frac{1}{2}}\in(0,1)$. Now, we can estimate the second term in (3.28) using the above inequality as follows
$\frac{np+2}{2-(k-1)n}\int_{\Omega}{{u}}^{p}$ $\leq\frac{d_{1}(p-1)}{p}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}+C_{6}.$ (3.29)
Again, using the Gagliardo-Nirenberg inequity, to get
$\int_{\Omega}{u}^{\frac{np+2}{n}}=$ $\left\|{u}^{\frac{p}{2}}\right\|^{\frac{2(np+2)}{np}}_{\frac{2(np+2)}{np}}\leq C_{7}\left(\left\|\nabla{u}^{\frac{p}{2}}\right\|^{\frac{np}{np+2}}_{{\mathcal{L}^{2}}(\Omega)}\,\,\left\|{u}^{\frac{p}{2}}\right\|^{\frac{2}{np+2}}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}+\left\|{u}^{\frac{p}{2}}\right\|_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}\right)^{\frac{2(np+2)}{np}}$
$\leq$ $C_{7}\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}\,\,\left\|{u}^{\frac{p}{2}}\right\|^{\frac{4}{np}}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}+C_{7}\left\|{u}^{\frac{p}{2}}\right\|^{\frac{2(np+2)}{np}}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}$
$\leq$ $C_{7}\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}\>\Big{\|}{u}\Big{\|}^{\frac{2}{n}}_{{\mathcal{L}^{1}}(\Omega)}+C_{7}\Big{\|}{u}\Big{\|}^{\frac{np+2}{n}}_{{\mathcal{L}^{1}}(\Omega)}$
$\leq$ $C_{8}\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{9}$
$\leq$ $C_{8}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}+C_{9}.$
Now, we can estimate the third term in (3.28) using the above inequality as follows
$\frac{-d_{1}(p-1)}{p}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}\leq\frac{-d_{1}(p-1)}{pC_{8}}\int_{\Omega}{u}^{\frac{np+2}{n}}+C_{10}.$ (3.30)
Substituting (3.29) and (3.30) in (3.28), we arrive at
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}+\frac{np+2}{2-(k-1)n}\int_{\Omega}{{u}}^{p}\leq$ $\frac{-d_{1}(p-1)}{pC_{8}}\int_{\Omega}{u}^{\frac{np+2}{n}}+2\epsilon\int_{\Omega}{u}^{\frac{np+2}{n}}$
$+C_{1}\int_{\Omega}{|\Delta{v}|}^{\frac{np+2}{2-(k-1)n}}+C_{2}\int_{\Omega}{|\Delta{w}|}^{\frac{np+2}{2-(k-1)n}}+C_{11}.$ (3.31)
Applying the variation-of-constants formula to (3.31), we obtain the following estimate with $\epsilon=\frac{d_{1}(p-1)}{4pC_{8}}$
$\int_{\Omega}{{u}}^{p}\leq$ $\frac{-d_{1}(p-1)}{2pC_{8}}\int^{t}_{s_{0}}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}\int_{\Omega}{u}^{\frac{np+2}{n}}+C_{1}\int^{t}_{s_{0}}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}\int_{\Omega}{|\Delta{v}|}^{\frac{np+2}{2-(k-1)n}}$
$+C_{2}\int^{t}_{s_{0}}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}\int_{\Omega}{|\Delta{w}|}^{\frac{np+2}{2-(k-1)n}}+C_{12},$ (3.32)
where $C_{12}=\int_{\Omega}{u}(\cdot,s_{0})+C_{11}\frac{2-(k-1)n}{np+2}$. From the Lemma 2.3, there exists $C_{p}>0$ such that
$C_{1}\int^{t}_{s_{0}}\int_{\Omega}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}{|\Delta{v}|}^{\frac{np+2}{2-(k-1)n}}\leq$ $C_{13}\int^{t}_{s_{0}}\int_{\Omega}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}|{u}|^{\frac{(np+2)l}{2-(k-1)n}}$
$+C_{14}\Big{\|}{v}(\cdot,s_{0})\Big{\|}^{\frac{np+2}{2-(k-1)n}}_{{\mathcal{W}}^{2,\frac{np+2}{2-(k-1)n}}}.$ (3.33)
By the similar manner, we have
$C_{2}\int^{t}_{s_{0}}\int_{\Omega}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}{|\Delta{w}|}^{\frac{np+2}{2-(k-1)n}}\leq$ $C_{15}\int^{t}_{s_{0}}\int_{\Omega}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}|{u}|^{\frac{(np+2)l}{2-(k-1)n}}$
$+C_{16}\Big{\|}{w}(\cdot,s_{0})\Big{\|}^{\frac{np+2}{2-(k-1)n}}_{{\mathcal{W}}^{2,\frac{np+2}{2-(k-1)n}}}.$ (3.34)
Substituting (3.33) and (3.34) in (3.32), we obtain that
$\int_{\Omega}{{u}}^{p}\leq$ $\frac{-d_{1}(p-1)}{2pC_{8}}\int^{t}_{s_{0}}\int_{\Omega}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}{u}^{\frac{np+2}{n}}$
$+C_{17}\int^{t}_{s_{0}}\int_{\Omega}e^{-\left(\frac{np+2}{2-(k-1)n}\right)(t-s)}{u}^{\frac{(np+2)l}{2-(k-1)n}}+C_{18},$ (3.35)
where $C_{18}=C_{12}+C_{14}\Big{\|}{v}(\cdot,s_{0})\Big{\|}^{\frac{np+2}{2-(k-1)n}}_{{\mathcal{W}}^{2,\frac{np+2}{2-(k-1)n}}}+C_{16}\Big{\|}{w}(\cdot,s_{0})\Big{\|}^{\frac{np+2}{2-(k-1)n}}_{{\mathcal{W}}^{2,\frac{np+2}{2-(k-1)n}}}$. Using the Young’s inequality with $l+k<1+\frac{2}{n}$, we have
$C_{17}\int_{\Omega}{u}^{\frac{(np+2)l}{2-(k-1)n}}\leq\frac{d_{1}(p-1)}{2pC_{8}}\int_{\Omega}{u}^{\frac{np+2}{n}}+C_{19},$ (3.36)
where $C_{19}=\left(\frac{d_{1}(p-1)}{2pC_{8}}\frac{\frac{2}{n}+1-k}{l}\right)^{\frac{-l}{\frac{2}{n}+1-(k+l)}}\frac{\frac{2}{n}+1-(k+l)}{\frac{2}{n}+1-k}C_{17}^{\frac{\frac{2}{n}+1-k}{\frac{2}{n}+1-(k+l)}}|\Omega|$. Substituting (3.36) in (3.35), finally, we arrive at
$\int_{\Omega}{{u}}^{p}\leq C_{20},\qquad\qquad\forall t\in(s_{0},T_{\mathrm{max}}).$
where the constant $C_{20}>0$.
Now, for $M_{0}\geq 1$, from Lemma 2.2 we have $1\leq M(t)\leq M_{0}$. Then (3.24) can be written as
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}\leq$ $-d_{1}p(p-1)\int_{\Omega}{{u}}^{p-2}|{\nabla{u}}|^{2}+\chi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{v}}$
$-\xi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{w}}.$
Applying the same procedure, one can obtain the boundedness. This completes the proof. ∎
Proof of Theorem 1.1. We can argue the proof by contradiction. Suppose that $T_{\mathrm{max}}<\infty$, utilizing the Lemma 2.4, the bounds (2.9), (2.10) and (2.16) holds for $t\in(s_{0},T_{\mathrm{max}})$. The blow-up criterion (2.1) contradicts with the above result that is, the maximum time is not finite. Finally, we done the following bounds with (3.4)
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq C_{17},\qquad\big{\|}{v}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\leq C_{18}\qquad\mathrm{and}\qquad\big{\|}{w}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\leq C_{19},$
for all $t>0$. This completes the proof.
4 Numerical Examples
This section discusses the results of the numerical simulations that clarify the behaviour of the solutions of the attraction-repulsion chemotaxis system in two space dimensions based on the finite difference method. In all of the examples below, we consider $d_{1}=d_{2}=d_{3}=\chi=\xi=\alpha=\beta=1$ and use a uniform grid $\Delta x=\Delta y=\frac{1}{201}$ in a square domain $[-0.5,0.5]\times[-0.5,0.5]$. Moreover, in the following examples, $\|u-u^{*}\|_{\infty}$ represents the maximum of difference between present time and previous time of $u(x,y,t)$.
Example 4.1 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})}\qquad\mbox{and}\qquad v(x,y,0)=w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
The attraction-repulsion chemotaxis system is tested in this example with and without a nonlocal source under same initial conditions for both $v$ and $w$. First we compute the cell densities for $\mu=0$ at a large time $t=10$ which are plotted in Figure 1.
((a))
((b))
((c))
Figure 1: Cell and chemical concentrations for $\mu=0$ at $t=10$.
Now compute the cell densities for $\mu=1$ for a large time $t=10$ and the obtained cell densities plotted in Figure 2.
((a))
((b))
((c))
Figure 2: Cell and chemical concentrations for $\mu=1$ at $t=10$.
From the above two situations, we observe that the solution decays in time and behaves like the fundamental solution of a heat equation. This is because the repulsion nullifies the attraction due to same initial conditions.
From the above two tests, we conclude that the nonlocal source term substantially reduces the convergence range of the solution and this source term may help to delay the blowup time. Moreover from the above discussions, we notice that when both the chemical concentrations $v$ and $w$ are initially the same, the solution cannot blowup. This numerical example particularly confirms the theoretical results of Jin[21].
Example 4.2 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=500e^{-50(x^{2}+y^{2})},\quad w(x,y,0)=250e^{-25(x^{2}+y^{2})}.$
Similar to the previous case, one can observe the convergence here. Hence it is concluded that whenever the attraction dominates the repulsion with an appropriate initial data, the solution converges for both the cases of $\mu=0$ and $1$.
((a))
((b))
((c))
Figure 3: Cell and chemical concentrations for $\mu=0$ at $t=10$.
((d))
((e))
((f))
Figure 4: Cell and chemical concentrations for $\mu=1$ at $t=10$.
Example 4.3 (Quick blowup of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=250e^{-25(x^{2}+y^{2})},\quad w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
This example deals with the behavior of the solution when the concentration of $w$ is initially higher than that of $v$. Figure 6 represents the cell densities at $t=3\times 10^{-5}$ for $\mu=0$. Unfortunately the solution immediately blows up after the time $t=2.5\times 10^{-5}$ and the spiky shape of the cell concentration is collapsed. Additionally we test the same system with a nonlocal source, the captured cell densities are figured out at time $t=3\times 10^{-5}$ in Figure 6.
((a))
((b))
Figure 5: Behaviour of $u$ for $\mu=0$ at $t=3\times 10^{-5}$.
((c))
((d))
Figure 6: Behaviour of $u$ for $\mu=1$ at $t=3\times 10^{-5}$.
From the two cases, one can see that the cells proliferate uncontrollably in the domain since the repulsion dominates the attraction.
Example 4.4 (Convergence of solution for small initial data).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=100e^{-10(x^{2}+y^{2})},\quad v(x,y,0)=25e^{-2.5(x^{2}+y^{2})},\quad w(x,y,0)=50e^{-5(x^{2}+y^{2})}.$
This example deals with the behaviour of the system (1.6) when repulsion dominates the attraction under a small initial data. Contrary to the previous Example 4.3, the solution converges with an appropriate small initial data. From this, we conclude that whenever repulsion dominates the attraction, the small initial data help to attain the convergence; this can be seen in Figure 8.
((a))
((b))
((c))
((a))
((b))
((c))
Figure 8: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
Example 4.5 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1,m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})}\qquad\mbox{and}\qquad v(x,y,0)=w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
Compared to solutions in Example 4.1, here the solution rather converges in a very short time and the exponent $m$ plays a crucial role in this convergence. This can be seen in the following Figure 9.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 9: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
Example 4.6 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1$, $m=1.8$, $\mu=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=500e^{-50(x^{2}+y^{2})},\quad w(x,y,0)=250e^{-25(x^{2}+y^{2})}.$
Figure 13 shows the cell densities at $t=10^{-4}$. One can see from the solution at time $t=10^{-4}$ that the cell density has almost reached the blowup state. But fortunately at time $t=10^{-3}$, the solution start to converge and the maximum value of $u$ keeps decreasing compared to the previous time $t=10^{-4}$. Now while we go for the next time $t=10^{-2}$, the solution converges absolutely and the growth of the cell density reaches equilibrium.
((a))
((b))
((c))
Figure 10: Behaviour of $u$ at $t=10^{-4}$.
((d))
((e))
((f))
Figure 11: Behaviour of $u$ at $t=10^{-3}$.
((a))
((b))
((c))
Figure 12: Behaviour of $u$ at $t=10^{-2}$.
((a))
((b))
((c))
Figure 13: Cell and chemical concentrations at $t=10$.
This experiment shows that, for specific values of the exponents $k,l$ and $m$, even if the initial condition is large, the system can be prevented from the blowup.
Example 4.7 (Quick blowup of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=250e^{-25(x^{2}+y^{2})},\quad w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
Quite similar to Example 4.3, the following figures indicate the cell densities captured at $t=3\times 10^{-5}$. Note that the solution seems to converge but at much later time $t=3\times 10^{-5}$, the solution suddenly blows up and the spiky shape is collapsed.
((a))
((b))
Figure 14: Behaviour of $u$ for $\mu=0$ at $t=3\times 10^{-5}$.
((c))
((d))
Figure 15: Behaviour of $u$ for $\mu=1$ at $t=3\times 10^{-5}$.
From this example, it is observed that the solution still blows up for a higher value of $m$ due to the domination of attraction by repulsion.
Example 4.8 (Convergence of solution for small initial data).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=100e^{-10(x^{2}+y^{2})},\quad v(x,y,0)=25e^{-2.5(x^{2}+y^{2})}\quad\mbox{and}\quad w(x,y,0)=50e^{-5(x^{2}+y^{2})}.$
This example deals with the behaviour of the system (1.6) when repulsion dominates the attraction for particular values of the exponents $k,l$ and $m$ under a small initial data. Unlike the previous Example 4.7, the solution converges with a small initial data.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 16: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
The system (1.6) admits imaginary solutions for values of $k$ greater than 1 and the captured cell densities are depicted in the examples below for various scenarios.
Example 4.9 (Imaginary solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.4$, $k=1.5$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\>v(x,y,0)=500e^{-50(x^{2}+y^{2})},\>w(x,y,0)=250e^{-25(x^{2}+y^{2})}.$
The behavior of the solution when $v>w$ initially, is illustrated in the figures below. It can be observed that for $\mu=0$, the solution transits to an imaginary state after $t=0.76\times 10^{-6}$, while for $\mu=1$, this transition occurs after $t=0.77\times 10^{-6}$.
((a))
((b))
Figure 17: Behaviour of $u$ for $\mu=0$ at $t=0.76\times 10^{-6}$.
((c))
((d))
Figure 18: Behaviour of $u$ for $\mu=1$ at $t=0.77\times 10^{-6}$.
Example 4.10 (Imaginary solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.4$, $k=1.5$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\>v(x,y,0)=250e^{-25(x^{2}+y^{2})},\>w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
The subsequent figures illustrate how the solution behaves when initial values $v<w$. In this scenario, it is observed that for $\mu=0$, the solution reaches an imaginary state after $t=0.74\times 10^{-6}$ and for $\mu=1$, this transitions occurs after $t=0.75\times 10^{-6}$.
((a))
((b))
Figure 19: Behaviour of $u$ for $\mu=0$ at $t=0.74\times 10^{-6}$.
((c))
((d))
Figure 20: Behaviour of $u$ for $\mu=1$ at $t=0.75\times 10^{-6}$.
Example 4.11 (Convergence of solution for small initial data).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.4$, $k=1.5$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=10e^{-10(x^{2}+y^{2})},\>v(x,y,0)=5e^{-5(x^{2}+y^{2})},\>w(x,y,0)=2.5e^{-2.5(x^{2}+y^{2})}.$
The behavior of the solution is illustrated in the following figures, assuming the initial values $v>w$. It is observed that with small initial data, the solution converges for $k=1.5$. These figures demonstrate the long-term behavior of the system (1.6) at $t=10$ for $\mu=0$ and $\mu=1$.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 21: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
Remark 4.1.
If $k>1$, using our approach as in the above Example 4.11, the convergence can be achieved for the cases $v(x,y,0)=w(x,y,0)$ and $v(x,y,0)<w(x,y,0)$.
5 Conclusion
This work analyzes the existence of classical solution to the attraction repulsion chemotaxis system with a nonlocal source and sublinear production. For $n\geq 2$, it is proved that the system (1.6) has a bounded classical solution when the parameters $k,l$ and $m$ satisfy some suitable conditions.
Next we have presented some comprehensive numerical examples that interpret the theoretical results. From the above graphical observations, we notice the following results.
  • Suppose the concentration of the attraction and repulsion chemicals are initially the same, the system (1.6) always converges.
  • When the attraction dominates the repulsion, the system (1.6) still converges.
  • On the other hand, when repulsion dominates the attraction, the solution suddenly blows up. In this case, preventing the blowup is very critical if the initial condition is large.
  • The system (1.6) has imaginary solutions for values of $k$ greater than 1. Nevertheless, the convergence can be achieved with small initial data.
Acknowledgments
GS thanks to Anusandhan National Research Foundation (ANRF), formerly Science and Engineering Research Board (SERB), Govt. of India for their support through Core Research Grant $(CRG/2023/001483)$ for this work.